OpenStudy (tyteen4a03):

log(10^2x - 56) - x = 0 Not sure how to tackle this question... Any pointers?

OpenStudy (anonymous):

what's the base 10 or e

OpenStudy (tyteen4a03):

@bhaskarbabu Base 10.

ganeshie8 (ganeshie8):

entire (2x-56) is in exponent ?

OpenStudy (tyteen4a03):

@ganeshie8 Nope, 10^2x and -56.

ganeshie8 (ganeshie8):

thats why u posted this :s

OpenStudy (tyteen4a03):

@ganeshie8 That's why I post what?

ganeshie8 (ganeshie8):

nvm :)

ganeshie8 (ganeshie8):

start with converting log to exponent

ganeshie8 (ganeshie8):

log(10^2x - 56) - x = 0 10^2x - 56 = 10^x

ganeshie8 (ganeshie8):

and let 10^x = k, and solve the quadratic maybe

OpenStudy (anonymous):

log(10^2x - 56) = x 10^2x - 56 = 10^x 10^2x -10^x - 56 =0 let 10^x=k k^2-k-56 = 0 solve for k

OpenStudy (tyteen4a03):

@bhaskarbabu Ah, I see. Thanks!